A back yard has a total length of 46 ft. And it is sectioned off into 3 parts. The first part is 2ft less than the second part and the third part is double the second part. Find the length of the first part.

Thank you, for your help

Answer by Cromlix(4381)

Second part = x
First part = x - 2
Third part = 2x
(x - 2) + x + 2x = 46
4x - 2 = 46
4x = 48
x = 12
So, First part = 10ft
Second part = 12ft
Third part = 24ft
Total = 46ft
Hope this helps.
:-)